# The diagnostic temperature discrepancy as evidence for non-Maxwellian coronal electrons

### 06 Mar 2026 v1

@@ Abstract
Two independent electron temperature diagnostics applied to the quiet solar corona yield systematically different results. Radio brightness temperatures from the Nançay Radioheliograph indicate $T_{\rm e}\approx 0.6$ MK, while hydrostatic scale-height modeling of the same plasma requires $T_{\rm e}\approx 1.5$ MK (Mercier & Chambe, 2015). Both diagnostics probe electrons; they disagree by a factor of $R=2.4\pm 0.3$. This discrepancy persists across an eight-year dataset spanning solar minimum and is consistent with LOFAR observations at lower frequencies (Vocks et al., 2018). We consider the propagation alternative (turbulent scattering of radio emission), which operates in the correct direction to suppress the apparent brightness temperature, but the ratio $R$ is invariant over the solar cycle despite expected variations in turbulence levels. We propose that the residual, cycle-invariant discrepancy reflects non-Maxwellian electron velocity distributions. Radio bremsstrahlung samples the distribution core (Chiuderi & Chiuderi Drago, 2004; Fleishman & Kuznetsov, 2014), while ionization rates and scale heights are dominated by the suprathermal tail (Owocki & Scudder, 1983). For kappa distributions, the predicted ratio is $\kappa/(\kappa-3/2)$; the observed $R=2.4$ implies $\kappa\approx 2$–3. This is consistent with spectroscopic measurements in active regions (Dudík et al., 2015) but in tension with perturbative theoretical predictions of $\kappa\approx 10$–25 (Cranmer, 2014). We make falsifiable predictions: Active Region cores should show a collapsed ratio ($R\leq 1.5$) as collisionality restores thermal equilibrium. Applying fluid transport equations (Spitzer-Härm conductivity) to plasmas with $\kappa\approx 2$–3 is physically invalid, but we do not compute the resulting heat flux, which remains an open problem.
@@ 3.1 Turbulent scattering
Radio waves propagating through a turbulent corona undergo scattering by electron density fluctuations. Scattering increases the apparent source size without changing the total flux, which strictly reduces the observed brightness temperature. This is a real effect. The question is whether it accounts for the full factor-of-2.4 discrepancy.
@@
Two independent constraints limit its role.
Geometric insufficiency. Scattering conserves total flux while increasing the apparent source solid angle: $T_{b,\rm app}=T_{b,\rm int}\times(\Omega_{\rm int}/\Omega_{\rm app})$. To suppress $T_{b}$ from 1.5 MK to 0.6 MK purely through scattering, the apparent source area must inflate by a factor of 2.5 (i.e., $\sim$150% broadening). Sharma & Oberoi (2020) directly measured the quiet Sun angular broadening at 25–30% in area. Even taking the upper end of their measured $\delta n/n$ range, the observed scattering is quantitatively insufficient by a factor of $\sim$5. This is a direct observational constraint, independent of any assumptions about cycle variability or turbulence models.
Solar cycle invariance. The ratio $R$ is stable over Mercier & Chambe’s eight-year dataset, spanning both solar maximum and the deep 2008–2009 minimum (the quietest Sun in a century). Coronal turbulence levels vary with solar activity: density fluctuation amplitudes, magnetic complexity, and the occurrence of transient events all increase toward solar maximum. If scattering drove the $T_{b}$ suppression, $R$ should track these variations. It does not. The stability of $R$ across a full solar cycle, combined with confirmation at different frequencies by LOFAR (Vocks et al., 2018), points to a cycle-invariant property of the electron distribution rather than a propagation effect.
Together, these constraints establish that scattering contributes to $T_{b}$ suppression but cannot account for the full discrepancy. The measured angular broadening can explain at most $\sim$20% of the required suppression. The remaining factor of $\sim$2 reflects an intrinsic property of the electron distribution.
@@ 6.1 Active Region core collapse
In Active Region cores ($n_{e}\geq 10^{9}$ cm-3), collisionality increases. The collision frequency scales as $\nu\propto n_{e}T^{-3/2}$. The thermalization timescale becomes shorter than the tail-regeneration timescale. The distribution relaxes toward Maxwellian: $\kappa\to\infty$, $R\to 1$.
Prediction: Active Region cores should exhibit $R\leq 1.5$, corresponding to $\kappa\geq 5$.
Falsification: If the diagnostic ratio in AR cores remains $\geq 2.2$ (indistinguishable from quiet Sun), this framework fails.
This prediction applies specifically to AR cores with quasi-steady heating, the one regime where neither passive filtration nor active driving operates efficiently. It does not apply to dynamic AR loops or transient features, where impulsive energy release can maintain low $\kappa$ despite high densities. In AR cores, gyroresonance emission at low harmonics of the electron cyclotron frequency can dominate over bremsstrahlung, complicating the interpretation of radio brightness temperatures. The test therefore requires either frequencies above the gyroresonance layers or careful isolation of the free-free component.
@@ 8.1 Summary
The diagnostic temperature discrepancy in the quiet solar corona ($R=T_{H}/T_{b}\approx 2.4$) is not a calibration artifact. It is built into the frequency-dependent spectral shape of the radio emission (Mercier & Chambe, 2015), stable over an eight-year solar cycle, and consistent with LOFAR observations at lower frequencies (Vocks et al., 2018). Turbulent scattering contributes to $T_{b}$ suppression but is quantitatively insufficient (the measured angular broadening accounts for at most $\sim$20% of the required suppression), and the ratio is invariant over the solar cycle (Section 3.1). Ion-electron temperature separation fails the solar cycle test (Section 3.2).
@@ 8.2 Relation to previous work
Chiuderi & Chiuderi Drago (2004) established the theoretical framework: kappa-distributed bremsstrahlung reduces quiet Sun radio brightness temperature relative to Maxwellian. Their work predates the Mercier & Chambe measurement and focused on microwave/cm-mm wavelengths without framing the result as a diagnostic ratio or extracting a specific $\kappa$. Our contribution is the inversion: applying their physics to Mercier & Chambe’s specific data to extract $\kappa$ from the observed discrepancy.
@@ 8.3 Open questions
We identify four questions that this paper raises but does not answer:
@@
    Scattering decomposition. What fraction of the observed $T_{b}$ suppression is due to scattering versus intrinsic non-Maxwellian effects? This requires simultaneous radio brightness, angular broadening, and spectroscopic constraints.

### 07 Apr 2026 v2

@@ Abstract
Two independent electron temperature diagnostics applied to the quiet solar corona yield systematically different results. Radio brightness temperatures from the Nançay Radioheliograph indicate $T_{\rm e}\approx 0.6$ MK, while hydrostatic scale-height modeling of the same plasma requires $T_{\rm e}\approx 1.5$ MK (Mercier & Chambe, 2015). Both diagnostics probe electrons; they disagree by a factor of $R=2.4\pm 0.3$. This discrepancy persists across an eight-year dataset spanning solar minimum and is consistent with LOFAR observations at lower frequencies (Vocks et al., 2018). We consider the propagation alternative (turbulent scattering of radio emission), which operates in the correct direction to suppress the apparent brightness temperature. Comparison with the FORWARD/PSIMAS Maxwellian model (Sharma & Oberoi, 2020) shows that the standard thermal structure predicts $\sim$1.6 MK; scattering accounts for the modest reduction toward observed MWA values but not the gap to 620 kK. The ratio $R$ is also invariant over the solar cycle despite measured variations in turbulence levels (Gautam et al., 2024). We propose that the residual discrepancy reflects non-Maxwellian electron velocity distributions. Radio bremsstrahlung samples the distribution core (Chiuderi & Chiuderi Drago, 2004; Fleishman & Kuznetsov, 2014), while ionization rates and scale heights are dominated by the suprathermal tail (Owocki & Scudder, 1983). For kappa distributions, the predicted ratio is $\kappa/(\kappa-3/2)$; the observed $R=2.4$ implies $\kappa\approx 2$–3. This is consistent with spectroscopic measurements in active regions (Dudík et al., 2015) but in tension with perturbative theoretical predictions of $\kappa\approx 10$–25 (Cranmer, 2014). We make falsifiable predictions: Active Region cores should show a collapsed ratio ($R\leq 1.5$) as collisionality restores thermal equilibrium. Applying fluid transport equations (Spitzer-Härm conductivity) to plasmas with $\kappa\approx 2$–3 is physically invalid, but we do not compute the resulting heat flux, which remains an open problem.
@@ 3.1 Turbulent scattering
Radio waves propagating through a turbulent corona undergo scattering by electron density fluctuations. Scattering increases the apparent source size and, when the medium has non-zero optical depth, absorption along the lengthened ray paths can reduce the total emergent flux (Thejappa & MacDowall, 2008). Both effects suppress the observed brightness temperature. This is a real effect. The question is whether it accounts for the full factor-of-2.4 discrepancy.
@@
Three independent constraints limit its role.
The FORWARD comparison. Sharma & Oberoi (2020) compared MWA observations with synthetic brightness temperature maps generated using the FORWARD package (Gibson et al., 2016), which computes Maxwellian thermal bremsstrahlung from the PSIMAS MHD coronal model. The FORWARD maps include the full coronal thermal structure (density, temperature, and magnetic field from data-driven MHD) but do not include scattering by density fluctuations, nor do they assume non-Maxwellian electron distributions. At 217 MHz near disk center, FORWARD predicts a peak $T_{b}$ of 1.61 MK; the MWA observes 1.50 MK (their Figure 4). At 162 MHz the gap widens to $\sim$25% (1.60 vs. 1.20 MK), and at 108 MHz to $\sim$42% (1.59 vs. 0.92 MK). The growing discrepancy at lower frequencies is consistent with scattering increasing at longer wavelengths, and the difference between FORWARD and MWA quantifies the scattering contribution.
The key point is that even at 108 MHz, where scattering produces its largest effect in the Sharma & Oberoi dataset, the MWA peak $T_{b}$ is 0.92 MK — still 50% above the 620 kK optically thick saturation measured by Mercier & Chambe (2015). Importantly, these are measurements in different optical depth regimes. The MWA observations at 108–217 MHz sample a regime where optical depth effects, thermal structure, and scattering all modulate $T_{b}$. Mercier & Chambe’s optically thick saturation below $\sim$200 MHz directly measures the electron temperature weighted by the emitting plasma. The standard Maxwellian model, with all thermal structure included, predicts $\sim$1.6 MK. Scattering accounts for the frequency-dependent reduction from 1.6 MK toward the MWA values. The gap between 1.6 MK and 620 kK is the discrepancy this paper interprets.
Geometric insufficiency. Purely flux-conserving scattering reduces $T_{b}$ by redistributing emission over a larger solid angle: $T_{b,\rm app}=T_{b,\rm int}\times(\Omega_{\rm int}/\Omega_{\rm app})$. To suppress $T_{b}$ from 1.5 MK to 0.6 MK through this mechanism alone, the apparent source area must inflate by a factor of 2.5 (i.e., $\sim$150% broadening). Sharma & Oberoi (2020) directly measured the quiet Sun angular broadening at 25–30% in area. Even taking the upper end of their measured $\delta n/n$ range, the observed broadening is quantitatively insufficient by a factor of $\sim$5. As noted above, scattering with absorption can reduce $T_{b}$ beyond the purely geometric limit, but the FORWARD comparison places a quantitative bound on the total scattering contribution across all mechanisms.
Solar cycle invariance. The ratio $R$ is stable over Mercier & Chambe’s eight-year dataset, spanning both solar maximum and the deep 2008–2009 minimum (the quietest Sun in a century). Coronal turbulence levels vary with solar activity: Gautam et al. (2024) measured turbulence cascade rates from WIND spacecraft data spanning 1995–2020 and found that the cascade rate is largest during solar maximum and smallest during solar minimum. Baseline quiet-Sun turbulence persists even at solar minimum, and some coronal heating models are based on MHD turbulence. But if scattering drove the bulk of the $T_{b}$ suppression, $R$ should track variations in turbulence levels across the cycle. It does not. The stability of $R$, including through the deepest minimum in a century, combined with confirmation at different frequencies by LOFAR (Vocks et al., 2018), points to a cycle-invariant property of the electron distribution rather than a propagation effect.
Together, these constraints establish that scattering contributes to $T_{b}$ suppression but cannot account for the full discrepancy. The FORWARD comparison shows that the standard Maxwellian model with realistic thermal structure starts at $\sim$1.6 MK; scattering reduces this toward the MWA values but leaves the gap to 620 kK unexplained. The remaining factor of $\sim$2 reflects an intrinsic property of the electron distribution.
@@ 6.1 Active Region core collapse
In Active Region cores ($n_{e}\geq 10^{9}$ cm-3), collisionality increases. The average collision frequency scales as $\nu\propto n_{e}T^{-3/2}$, but for individual electrons the rate scales as $v^{-3}$ (Shoub, 1983): low-velocity (thermal core) electrons thermalize far faster than high-velocity (suprathermal tail) electrons. The bulk of the distribution therefore relaxes toward Maxwellian while residual suprathermal tails persist on longer timescales. The resulting distribution is no longer a clean kappa — it is closer to a Maxwellian core plus a residual power-law tail, analogous to the solar wind core+halo decomposition.
In this partially thermalized regime, each diagnostic responds differently. Radio bremsstrahlung still samples the bulk, but the bulk has thermalized upward — $T_{\rm core}$ has relaxed toward $T_{\rm eff}$ — so $T_{b}$ increases relative to the quiet Sun case. The scale height and ionization balance remain sensitive to the tail (if it persists), so $T_{H}$ stays approximately constant. $R$ therefore decreases because $T_{b}$ has increased while $T_{H}$ has not, but $R$ does not reach unity as long as a residual tail keeps $T_{H}$ above $T_{b}$.
Prediction: $R$ should decrease systematically with increasing density, tracing the transition from tail-dominated to collision-dominated regimes. Different AR structures at different densities — quiet Sun, AR periphery, AR core — should trace out a continuous $R(n_{e})$ curve that maps the competition between tail regeneration and collisional thermalization. In AR cores with quasi-steady heating (Antiochos et al., 2003), $R$ should fall well below the quiet Sun value of 2.4.
Falsification: If the diagnostic ratio in AR cores remains $\geq 2.2$ (indistinguishable from quiet Sun), this framework fails, because even partial thermalization should produce measurable collapse.
The kappa inversion formula $\kappa=3R/[2(R-1)]$ assumes a kappa distribution. In the partially thermalized regime, the distribution is a hybrid. The formula still yields a number, but that number is an effective $\kappa$ characterizing the diagnostic ratio rather than the distribution shape. It remains useful as a single-parameter description of how non-Maxwellian the plasma appears to the core-versus-tail comparison, but it does not map cleanly onto the distribution function in this intermediate regime.
This prediction applies specifically to AR cores with quasi-steady heating — “quasi-steady” referring to the time-averaged heating rate being approximately constant (Antiochos et al., 2003), not the absence of dynamic activity. Small-scale reconnection events are ubiquitous, but the time-averaged collisionality they produce determines whether the tail regeneration rate can sustain a non-Maxwellian distribution. The prediction does not apply to dynamic AR loops or transient features, where impulsive energy release can maintain low $\kappa$ despite high densities. In AR cores, gyroresonance emission at low harmonics of the electron cyclotron frequency can dominate over bremsstrahlung, complicating the interpretation of radio brightness temperatures. The test therefore requires either frequencies above the gyroresonance layers or careful isolation of the free-free component.
@@ 8.1 Summary
The diagnostic temperature discrepancy in the quiet solar corona ($R=T_{H}/T_{b}\approx 2.4$) is not a calibration artifact. It is built into the frequency-dependent spectral shape of the radio emission (Mercier & Chambe, 2015), stable over an eight-year solar cycle, and consistent with LOFAR observations at lower frequencies (Vocks et al., 2018). Turbulent scattering contributes to $T_{b}$ suppression but cannot account for the full discrepancy: the standard Maxwellian model with realistic thermal structure predicts $\sim$1.6 MK (Sharma & Oberoi, 2020; Gibson et al., 2016); scattering reduces this toward observed MWA values but leaves the gap to 620 kK unexplained (Section 3.1). Ion-electron temperature separation fails the solar cycle test (Section 3.2).
@@ 8.2 Relation to previous work
Chiuderi & Chiuderi Drago (2004) established the theoretical framework: kappa-distributed bremsstrahlung reduces quiet Sun radio brightness temperature relative to Maxwellian. Their work predates the Mercier & Chambe measurement and focused on microwave/cm-$m$ wavelengths without framing the result as a diagnostic ratio or extracting a specific $\kappa$. Our contribution is the inversion: applying their physics to Mercier & Chambe’s specific data to extract $\kappa$ from the observed discrepancy.
@@ 8.3 Open questions
We identify five questions that this paper raises but does not answer:
@@
    Scattering decomposition. The FORWARD comparison (Section 3.1) provides a first-order decomposition of scattering and intrinsic effects at 108–217 MHz. Extending this into the optically thick regime below $\sim$200 MHz, where the diagnostic discrepancy is measured, requires forward-modeling with scattering included in the radiative transfer. Does the residual discrepancy persist when scattering is self-consistently incorporated?